OpenStudy (anonymous):

Solve for x:x^2-sq. root(15)x-2i=0

OpenStudy (zzr0ck3r):

\[x^2-\sqrt{15x}-2i=0?\]

OpenStudy (anonymous):

x is outside the root

OpenStudy (zzr0ck3r):

\[x^2-\sqrt{15}x-2i=0?\]

OpenStudy (anonymous):

yup

OpenStudy (zzr0ck3r):

\[x^2-\sqrt{15}x-2i=0\\(x-\frac{\sqrt{15}}{2})^2= 2i+(\frac{\sqrt{15}}{2})^2\\x=\pm\sqrt{2i+\frac{14}{4}}+\frac{\sqrt{15}}{2}\]

OpenStudy (zzr0ck3r):

do you now how to complete the square?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

that should say 15/4 not 14/4

OpenStudy (anonymous):

k

OpenStudy (zzr0ck3r):

ok, well that's all I did

OpenStudy (zzr0ck3r):

\[ax^2+bx+c=0\\a(x+\frac{b}{2})^2=-c+a(\frac{b}{2})^2\]

OpenStudy (anonymous):

got it

OpenStudy (zzr0ck3r):

good deal